OpenStudy (uri):

Halp meh pliss, The number of positive values of x less than 360 degrees which satisfy the equation sin 1/2x =cos x is? a)0 b)1) c)2 d)3 e)4

OpenStudy (saifoo.khan):

You do math too? D:

OpenStudy (hba):

Mujhe bhi yaqeen nahi horaha :P

OpenStudy (anonymous):

looks like there are two http://www.wolframalpha.com/input/?i=sin%28x%2F2%29%2Ccos%28x%29+domain+0..2pi

OpenStudy (yrelhan4):

write cosx in terms of sin^2 x/2 and solve!

OpenStudy (uri):

1/2x=sin^2x/2 Right ?

OpenStudy (yrelhan4):

cosx= 1-2sin^2 (x/2)

OpenStudy (yrelhan4):

you get a quadratic.. solve for x..

OpenStudy (uri):

2sin^2(x/2)+cosx-1=0

OpenStudy (yrelhan4):

2t^2 + t -1=0 where t=sin(x/2)

OpenStudy (uri):

Quad,breaking the middle term or completing the square method?

OpenStudy (yrelhan4):

breaking the middle term.. you would get (2t-1)(t+1)=0 where t=sin(x/2)

OpenStudy (uri):

eh,2t-1=0=)t=1/2 t+1=0=)t=-1

OpenStudy (yrelhan4):

yup!

OpenStudy (uri):

Sin(x/2)=1/2 Sin(x/2)=-1

OpenStudy (yrelhan4):

so there would be 3 values of x/2 in 0 to 360.. 30,120,270. so x=60,240,540.. since you want from 0 to 360.. there would be 2 solns..

OpenStudy (uri):

it was Sin x/2

OpenStudy (yrelhan4):

if sin(x/2) = 1/2 , what is x? 0>x>360

OpenStudy (hba):

Oh so, sin x/2 =cos x ?

OpenStudy (yrelhan4):

1- sin^2(x/2)= cosx @hba

OpenStudy (uri):

em BOREEEEEED.

OpenStudy (yrelhan4):

look.. let x/2=k .. sin(k)=1/2 --> k=30, 120.. --> x=60, 240

OpenStudy (yrelhan4):

cosx= 1-2sin^2 (x/2) @hba missed the two..

OpenStudy (yrelhan4):

thik hai? @uri

OpenStudy (sirm3d):

@yrelhan4 since \(\displaystyle 0 \leq x \leq 360\) then \(\displaystyle 0 \leq \frac{x}{2} \leq 180\)

OpenStudy (sirm3d):

now, \(\sin x/2 = 1/2\) therefore \(x/2=30, 150\) or \(x=60, 300\)

OpenStudy (uri):

Yrelhan GUESSSS WHAT?! I donyt undastand aaaanythang nd ima sho confuzzeld tooh,I caant eveen speel corractly. o.O

OpenStudy (yrelhan4):

lol.. i am such a fool.. you are right! @sirm3d

OpenStudy (uri):

Wut iz the aaansweeer?! none of my option matches your answer!

OpenStudy (yrelhan4):

i just wrotw150 value wrong, right? why am i a fool then? :P

OpenStudy (sirm3d):

the other equation, \(\sin x/2 = -1\) has no solution in \(0 \leq \frac{x}{2} \leq 180\)

OpenStudy (yrelhan4):

@uri 2.. final!

OpenStudy (uri):

@sirm3d sinx/2=1 has no solution so we will leave it? xD

OpenStudy (yrelhan4):

@uri its sin(x/2) = -1

OpenStudy (uri):

k.

OpenStudy (sirm3d):

@uri, just count the number of solutions: \(x=60^0,\;300^0\)

OpenStudy (uri):

@sirm3d 2 solutions?

OpenStudy (sirm3d):

yup. (C) 2

OpenStudy (yrelhan4):

are kitni baar puchegi? @uri :P

OpenStudy (uri):

AMG THANKYOUUUUUUUUUUUUU!AMG THANKYOUU SOO MUCHHHHH!!!! :D

OpenStudy (yrelhan4):

mera medal.. :'( .. aadha to maine solve kiya..

OpenStudy (yrelhan4):

huh!

OpenStudy (uri):

You have 2 now :P

OpenStudy (yrelhan4):

haha.. he's got 2 now too!

OpenStudy (uri):

Happy? :D

OpenStudy (yrelhan4):

ha! yes.. :P

OpenStudy (uri):

Yw xD